OpenStudy (anonymous):

for 4/7-1/2 can 7 and 2 go into 14

OpenStudy (anonymous):

Yep 7 and 2 both share a common denominator of 14 good job so let's start with 4/7 what do you have to multiply 7 by to get 14

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Good now multiply 7 by 2 to get the denominator and multiply the numerator which is 4 by 2 as well

OpenStudy (anonymous):

what does 4/7=?

OpenStudy (anonymous):

8/14

OpenStudy (anonymous):

Good job! Now lets work on the other fraction....

OpenStudy (anonymous):

1/2 what do you have to multiply 2 by to get 14?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

2/14

OpenStudy (anonymous):

Yep so multiply 2 by 7 to get 14 as your denominator and do the same thing to the top of the fraction what does 1/2=?

OpenStudy (anonymous):

You still have to multiply 2*7 to finish correcting the fraction

OpenStudy (anonymous):

wat

OpenStudy (anonymous):

so 1/2=7/14

OpenStudy (anonymous):

SO when I say do the same thing to the top of the fraction I mean multiply the top number of the fraction by the same number you multiplied by on the bottom of the fraction

OpenStudy (anonymous):

ok the answer was 1 14

OpenStudy (anonymous):

Your problem should now look like this: 8/14-7/14 subtract now and yep you got the answer right good job

OpenStudy (anonymous):

u plz not say good job then i feel like a kindergatiner

OpenStudy (anonymous):

ha ha sorry my bad

OpenStudy (anonymous):

if the denominaters is 5 or 6 is th lcm 11

OpenStudy (anonymous):

I think it would be 30....

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

cause lcm is The smallest (non-zero) number that is a multiple of two or more numbers.

OpenStudy (anonymous):

so 5*6=30 and there isn't any lower number that they both go into evenly other than 30

OpenStudy (anonymous):

i meant lcd

OpenStudy (anonymous):

oh ok lol no it would be 30 still for the same reason

OpenStudy (anonymous):

$!@#

OpenStudy (anonymous):

ha ha sorry

OpenStudy (anonymous):

ok so 30 then wat im confused on this one

OpenStudy (anonymous):

5*6=30 6*5=30 both 5 and 6 can't both go into anything that is lower than thirty

OpenStudy (anonymous):

oh ok the top numbers are 1/5 and 5/6

OpenStudy (anonymous):

Ok does the problem ask you to add or subtract them or what?

OpenStudy (anonymous):

1/5=6/30 5/6=25/30

OpenStudy (anonymous):

subtract

OpenStudy (anonymous):

no add

OpenStudy (anonymous):

ok so 6/30-25/30=-19/30

OpenStudy (anonymous):

ok so adding it would be 6/30+25/30=31/30 simplify 1(1/30)

OpenStudy (anonymous):

now if the denominaters were 5 and 2 couldnt the lcd be 10

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now if the denominaters were 5 and 2 couldnt the lcd be 10